Question 2.8 [2, 3, 3]



Suppose that a telemarketer has a 12% chance of making a sale on any given call. If the



telemarketer makes average of 5 calls per hour, calculate:



a) The probability that the telemarketer will make exactly 2 sales during the shift.



b) The probability that the telemarketer will makes more than 4 sales during the shift.



c) What is the probability that the telemarketer makes more than 2 call during two hours?

1
Expert's answer
2022-06-03T13:19:48-0400

Let "X=" the number of sales: "X\sim Bin(n, p)"

a)

"n=5, p=0.12, q=1-p=0.88"


"P(X=2)=\dbinom{5}{2}(0.12)^{2}(0.88)^{5-2}"

"=0.098131968"

b)

"n=5, p=0.12, q=1-p=0.88"


"P(X>4)=P(X=5)=\dbinom{5}{5}(0.12)^{5}(0.88)^{5-5}"

"=0.0000248832"

c)

"n=10, p=0.12, q=1-p=0.88"


"P(X>2)=1-P(X=0)-P(X=1)"

"-P(X=2)=1-\dbinom{10}{0}(0.12)^{0}(0.88)^{10-0}"

"-\dbinom{10}{1}(0.12)^{1}(0.88)^{10-1}-\dbinom{10}{2}(0.12)^{2}(0.88)^{10-2}"

"=0.10868179372"